Provethat12 +32 +52 +⋯+(2n+1)2 = (n+1)(2n+1)(2+ 3)∕3 whenever is a nonnegative integer.


1
Expert's answer
2021-08-31T12:55:57-0400

Let us prove that "1^2 +3^2 +5^2 +⋯+(2n+1)^2 = \frac{(n+1)(2n+1)(2n + 3)}{3}" using method of mathematical induction.


For "n=0" the left side is equal to "1^1=1," and the right side is equal to "\frac{(0+1)(2\cdot 0+1)(2\cdot 0 + 3)}{3}\frac{3}{3}=1."


Let for "n=k" the formula "1^2 +3^2 +5^2 +⋯+(2k+1)^2 = \frac{(k+1)(2k+1)(2k + 3)}{3}" is true.


Let us prove for "n=k+1:"


"1^2 +3^2 +5^2 +⋯+(2k+1)^2+(2(k+1)+1)^2 = \frac{(k+1)(2k+1)(2k + 3)}{3}+(2k+3)^2"


"= (2k+3)(\frac{(2k^2+3k+1}{3}+(2k+3))=(2k+3)\frac{2k^2+3k+1+6k+9}{3}=(2k+3)\frac{2k^2+9k+10}{3}"


"=(2k+3)\frac{(k+2)(2k+5)}{3}=\frac{(k+2)(2k+3)(2k+5)}{3}=\frac{((k+1)+1)(2(k+1)+1)(2(k+1)+3)}{3}."


We conclude that this formula is true for any non-negative integer number "n."